OpenStudy (anonymous):

Todd currently has $5,000. What was the value of his money four years ago if he's earned 4 percent interest each year?

OpenStudy (anonymous):

Kate wants to have $25,000 in 16 years. How much does she need to invest if the interest rate is 6 percent per year

OpenStudy (anonymous):

You didn't actually finish your question but from the multiple choice answers it would seem that you meant to say: "Interest rate is 6 percent per annum compounded annually" instead of the cryptic: "Interest rate is 6 percent per?" which you wrote. You will, or at least should remember that the future value equation is: FV = PV * ( 1 + i )^t Where FV is Future value PV is Present value i is the true effective annual interest rate as a fraction of 1 t is the number of years that the investment will be held for. You can rewrite this equation to read: PV = FV / ( 1 + i )^t Plug in your numbers and you would have: PV = $25,000 / ( 1 + 0.06 )^16 .: PV = $9,841.16